OpenStudy (anonymous):

How would you use the elimination method to solve the following system of equations. 2x + 9y = 4 3x + 7y = –7

OpenStudy (anonymous):

Would it just be like a normal equation. ? 2x + 9y = 4 3x + 7y = –7 5x+16y= -3 ??????????

OpenStudy (cwrw238):

no - u need to eliminate either the x term or y term by adding or subtracting the equations you cant do that as they stand because the coefficients ( that is number before the x and y are different)

OpenStudy (cwrw238):

one way to eliminate x is by multiplying first equation by 3 and second by -2 and then adding can you do that operation?

OpenStudy (cwrw238):

when multiplying equations be sure to multiply EACH TERM

OpenStudy (cwrw238):

are you with me?

OpenStudy (anonymous):

(2x + 9y = 4) times 3 (3x + 7y = –7) times 2 so every term must be timed.

OpenStudy (anonymous):

@cwrw238 sorry it took me so long. Okay i miltipled each equation. now do i add them or subtract?

OpenStudy (anonymous):

subtract.

OpenStudy (anonymous):

type what you got here.

OpenStudy (anonymous):

6x+27y=12 6x+14y=-14 13y=26 /13 /13 y=2 now how do I solve for x? put 2 into one of the equations?

OpenStudy (anonymous):

yep, that is correct.